I am confused. o.k. the problem is xsquared+7x+9=0 i do not get how to find the vertex . i know it has to do with the it being max or miminmum . right ?
Thanks for the help !!
- confused

Answer by stanbon(75887)

I am confused. o.k. the problem is x^2+7x+9=y i do not get how to find the vertex . i know it has to do with it being max or miminmum . right ?
---------------------
You have a quadratic with a=1 and b=7
The vertex occurs when x= -b/2a = -7/2
The y value when x=-7/2 = (-7/2)^2+7(-7/2)+9 = -13/4
---------------
The vertex is at (-7/2 , -13/4)
=================================
Cheers,
Stan H.